### v2

# Opinion Clustering under the Friedkin-Johnsen Model: Agreement in Disagreement

Abstract
The convergence of opinions in the Friedkin-Johnsen (FJ) framework is well studied, but the topological conditions leading to opinion clustering remain less explored. To bridge this gap, we examine the role of topology in the emergence of opinion clusters within the network. The key contribution of the paper lies in the introduction of the notion of topologically prominent agents, referred to as Locally Topologically Persuasive (LTP) agents. Interestingly, each LTP agent is associated with a unique set of (non-influential) agents in its vicinity. Using them, we present conditions to obtain opinion clusters in the FJ framework in any arbitrarily connected digraph. A key advantage of the proposed result is that the resulting opinion clusters are independent of the edge weights and the stubbornness of the agents. Finally, we demonstrate that by suitably placing LTP agents, one can design networks that achieve any desired opinion clustering.
{IEEEkeywords}
Opinion dynamics; Friedkin-Johnsen model; Clustering of opinions.
1 INTRODUCTION
The recent evolution of social media networks has generated a keen interest among researchers to study the impact of social interactions on public opinion. Several mathematical models such as the DeGroot’s model [1], the FJ model [2], the Hegselmann-Krause model [3], and the Biased assimilation model [4], explain complex aspects of human interactions such as individual biases and homophily. Among these, the FJ model is most analytically tractable and explains the frequently observed phenomenon of disagreement among closely interacting individuals [5]. Moreover, it is widely popular because it is empirically validated on real-world datasets [6], [7].
The opinions evolving under the FJ model converge to disagreement due to the presence of biased individuals, who are called stubborn agents. In [2], the authors show that if each non-stubborn agent in the network is reachable from a stubborn agent, the final opinions depend on the initial opinions of only the stubborn agents. Reference [8] introduces the notion of oblivious agents, who are non-stubborn and do not have a path from any stubborn agent. The algebraic and graph-theoretic conditions that ensure convergence in networks with oblivious agents are presented in [8] and [9], respectively.
In opinion dynamics, emergent behaviours such as consensus, polarisation and opinion clustering are of interest to explain and achieve a desired social outcome. Several works [6, 10, 11, 12] have examined the FJ framework to reach these behaviours. In [6], the authors design stubbornness values that lead to behaviours such as consensus and polarisation. On the other hand, reference [10] presents topology-based conditions for graphs with two stubborn agents that result in polarisation. Furthermore, the authors in [11] examine the FJ model with randomised interactions and reveal that the opinions of non-stubborn agents can exhibit oscillations in presence of stubborn agents. In contrast, the present work focuses on opinion clustering, which is less-explored in the FJ framework. Opinion clustering occurs when a group of two or more agents converge to the same final opinion. It represents swarm behaviour and has applications in task distribution [13] and formation control [14]. In the sociological framework, opinion clustering represents the formation of groups with aligned views.
In this work, we examine the network topology towards achieving opinion clustering. Recent works [10, 11] in the FJ framework highlight the impact of topology in shaping emergent behaviours. More recently, the study of network topology is receiving attention to analyse the effects of recommender algorithms on opinion formation. The seminal works [15] and [16] contributed significantly by quantifying the influence of a stubborn agent using hitting probabilities of random walks over the network. These studies enabled the design of topology-based interventions (recommendations) for desirably shaping opinions [16, 17]. In the present work, we further explore the relation between the network topology and final opinions and address the following problem: Under what topology-based conditions, does a set of agents converge to the same final opinion (form an opinion cluster)? We leverage this relationship to obtain desired opinion clustering. In [12], the authors present the graph-theoretic conditions to achieve opinion clustering in the FJ framework in networks with oblivious agents. However, the result applies only to networks where each cycle contains at most one stubborn agent. Contrary to [12], our work aims to explain the emergence of opinion clusters in the FJ framework in arbitrarily connected digraphs. With this, we now highlight our major contributions:
  • We define the notion of the Locally Topologically Persuasive (LTP) agent. Each LTP agent is associated with a unique set of non-influential agents in its vicinity.
  • We show that an LTP agent and the agents associated with it always form an opinion cluster under the FJ model. This result generalises the topology-based conditions presented in [12] to arbitrarily connected digraphs.
Paper Organisation: Sec.2 introduces the notations and the relevant preliminaries. Sec. 3 motivates the study of the formation of opinion clusters. Sec. 4, introduces the LTP agents and presents the conditions for opinion clustering. Finally, Sec. 5 concludes with insights into future directions.
2 Notations and Preliminaries
[…]
A walk is an ordered sequence of nodes such that each pair of consecutive nodes forms an edge in the graph. A cycle is a walk whose initial and final nodes coincide. A walk in which no node is repeated is a path. A path from node $x$ to node $y$ traverses node $z$ if the sequence of nodes that form this path contains $z$. A graph is aperiodic if the lengths of all its cycles are coprime. An undirected graph is connected if a path exists between every pair of nodes. A directed graph (digraph) is a strongly connected graph if a directed path exists between every pair of nodes in the graph. A digraph is weakly connected if the undirected version of the digraph is connected. A maximal subgraph of $\mathcal{G}$ which is strongly connected forms a strongly connected component (SCC). An SCC is an independent SCC (iSCC) if every node in the SCC has all its in-neighbours within the same SCC.
3 FJ model and emergent behaviours
Consider a network $\mathcal{G}$ of $n$ agents with vector $\mathbf{x}(k)=[x_{1}(k),...,x_{n}(k)]$ denoting the opinions of the agents at the $k^{th}$ instance. Under the FJ model, the opinions evolve as:
$\mathbf{x}(k+1)=(I-\beta)W\mathbf{x}(k)+\beta\mathbf{x}(0)$ (1)
[…]
Suppose $\mathcal{G}$ is weakly connected and there is a non-stubborn agent who does not have a directed path from any stubborn agent. Such agents are called oblivious agents [8]. Without loss of generality, we renumber the nodes in $\mathcal{G}$ such that the nodes $\{1,2,...,n_{o}\}$ are the oblivious agents and the rest are the non-oblivious, where $n_{o}\in[n]$. Additionally, we consider the oblivious agents that form an iSCC are grouped together. Now, we can re-write the FJ model (1) as:
$\mathbf{x}_{1}(k+1)$ $=W_{11}\mathbf{x}_{1}(k)$
$\mathbf{x}_{2}(k+1)$ $=(I-\bar{\beta})\big(W_{21}\mathbf{x}_{1}(k)+W_{22}\mathbf{x}_{2}(k)\big)+\bar{\beta}\mathbf{x}_{2}(0)$
where $\mathbf{x}_{1}(k)\in\mathbb{R}^{n_{o}}$ and $\mathbf{x}_{2}(k)\in\mathbb{R}^{(n-n_{o})}$ denote the opinions of oblivious and non-oblivious agents. The matrix $W$ gets partitioned as $W=\begin{bmatrix}W_{11}&0\\ W_{21}&W_{22}\end{bmatrix}$ and the diagonal matrix $\bar{\beta}$ denotes the stubbornness of non-oblivious agents. Throughout this article, we consider that $\mathcal{G}$ satisfies the following:
Assumption 1
Each iSCC in $\mathcal{G}$ composed of the oblivious agents with two or more agents is aperiodic.
Assumption 1 is the necessary and sufficient condition for convergence under the FJ model [9].
Lemma 1 ([8, 9])
Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents such that Assumption 1 holds. The opinions evolving under the FJ model (1) converge to,
$\mathbf{x}_{1}^{*}$ $=W_{11}^{*}\mathbf{x}_{1}(0)$
$\mathbf{x}_{2}^{*}$ $=(I-(I-\bar{\beta})W_{22})^{-1}\big((I-\bar{\beta})W_{21}W_{11}^{*}\mathbf{x}_{1}(0)$
$+\bar{\beta}\mathbf{x}_{2}(0)\big)$
where $\mathbf{x}^{*}=[\mathbf{x}_{1}^{*},\mathbf{x}_{2}^{*}]$ denotes the final opinions and $W_{11}^{*}=\lim_{k\to\infty}W_{11}^{k}$.
From Lemma 1, it clearly follows that the final opinions depend on the initial opinions of only the stubborn agents and the oblivious agents in iSCCs. Hence, these agents are referred to as the influential agents. The set of all influential agents in $\mathcal{G}$ is denoted by $\mathcal{I}$.
The reference [9] establishes that the final opinions lie in the convex hull of the initial opinions of influential agents. In [15], the coefficients of each of these convex combinations (that determine the final opinions) are related to the network topology (captured by $W$) and the stubbornness of agents. Depending on these factors, the final opinions of agents (except the oblivious agents in an iSCC) generally converge to distinct values, which is termed as disagreement.
Disagreement is a more prevalent in social networks as compared to global consensus [6]. A special case of disagreement occurs when a subgroup of two or more agents converges to the same final opinion and forms an opinion cluster. An opinion cluster is defined as:
Definition 1
(Opinion Cluster) A set of agents $\mathcal{C}\subset\mathcal{V}$ form an opinion cluster if for any initial conditions $\mathbf{x}(0)\in\mathbb{R}^{n}$
$x_{i}^{*}-x_{j}^{*}=0\quad\forall\ i,j\in\mathcal{C}$
where $x_{k}^{*}$ denotes the final opinion of agent $k\in\mathcal{V}$.
From a sociological perspective, the existence of opinion clusters captures a more realistic distribution of opinions in society, where local consensus occurs in groups (echo chambers) despite widespread disagreement. The formation of opinion clusters has been examined in opinion dynamics [3], however, the role of network topology leading in this phenomenon remains less explored. In [12], the authors present topology-based conditions for the formation of opinion clusters in digraphs where each cycle has at most one stubborn agent. However, these conditions are insufficient to explain the emergence of opinion clusters in the digraph in Example 1 that contains a cycle with multiple stubborn agents.
Example 1
Consider a strongly connected network $\mathcal{G}$ in Fig. 1(a) with stubborn agents $2$ and $6$. The cycle in $\mathcal{G}$ formed by agents $1,2$ and $6$ contains two stubborn agents. Let the stubbornness of the stubborn agents be $\beta_{2}=0.3$, and $\beta_{6}=0.6$. For initial opinions from a uniform distribution on $[0,5]$, the final opinions under the FJ model (1) form three opinion clusters as shown in Fig. 1(b).
(a) Network $\mathcal{G}$
(b) Evolution of opinions $\mathcal{G}$.
Figure 1: Formation of opinion clusters. Throughout this paper, we represent the stubborn and non-stubborn agents by red and orange nodes, respectively.
In this work, our objective is to establish conditions under which a subgroup of agents in any arbitrarily connected digraph forms an opinion cluster in the FJ framework. Using these conditions, we can use network interactions as a tool to form desired opinion clusters.
Remark 1
The agents within an opinion cluster converge to the same final opinions, regardless of initial opinions. This implies that each influential agent $s\in\mathcal{I}$ exerts an equal impact on all agents in the cluster. Thus, this work also presents the network structures under which a group of agents are equally affected by the influential agents.
4 Topology-based opinion clustering
Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (1). Let $q\in\mathcal{V}\setminus\mathcal{I}$ be a non-influential agent in $\mathcal{G}$. The final opinion $x_{q}^{*}$ depends on the initial opinions of only those influential agents that have a path to $q$. Let $\mathcal{I}_{q}\subseteq\mathcal{I}$ denote the set of influential agents that have a path to $q$. Certain topological properties in a network endow some agents with the special property of being ‘locally’ influential; we refer to them as LTP agents. More formally:
Definition 2
An agent $p\in\mathcal{V}$ is called an LTP agent if the following hold:
  1. (i)
    there exists an agent $q\in\mathcal{V}\setminus\mathcal{I}$ such that every path in $\mathcal{G}$ from each influential agent $s\in\mathcal{I}_{q}$ to $q$ (where $p\neq q$) traverses $p$,
  2. (ii)
    additionally, if $p\in\mathcal{V}\setminus\mathcal{I}$, then there does not exist an agent $c\in\mathcal{V}$ such that all paths from each agent $s\in\mathcal{I}_{p}$ to $p$ traverse $c$.
The non-influential agent $q$ is then said to be persuaded by $p$.
The set of all non-influential agents persuaded by an LTP agent $p$ is denoted by $\mathcal{N}_{p}$. If $p\in\mathcal{I}$ is an influential agent, condition (i) is sufficient to ensure that $p$ is an LTP agent. The condition (ii) in Defn. 2 is applicable only if $p$ is a non-influential agent. It conveys that a non-influential agent $p$, persuaded by an LTP agent $c$ (i.e $p\in\mathcal{N}_{c}$), is not an LTP agent itself. The following example illustrates which agents are LTP agents and which are the agents it persuades.
Example 2
Consider the network $\mathcal{G}$ shown in Fig. 1(a). Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential. Fig 2(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses $3$. Further, the paths from $6$ to $4$ (and $5$) also traverse $3$ as shown in Fig. 2(b). Since all paths from $2$ and $6$ to $3$ do not traverse any common node. Thus, $3$ is an LTP agent and $\mathcal{N}_{3}=\{4,5\}$. Similarly, $6$ is also an LTP agent and $\mathcal{N}_{6}=\{1\}$.
Note that each path from $2$ and $6$ to $5$ traverses $4$. Hence, $4$ satisfies condition 1) in Definition 2 However, since $4\in\mathcal{N}_{3}$, it is not an LTP agent.
(a) Paths from $2$ to $4$ (and $5$) traverse 33
(b) Paths from $6$ to $4$ (and $5$) traverse 33
Figure 2: Each path from the stubborn agents to 44 and 55 traverses $3$.
Remark 2
The notion of an LTP agent is analogous to that of cut vertices in undirected graphs (and strong articulation points in digraphs). A node $v$ is a cut vertex (or a strong articulation point in a strongly connected digraph) if and only if there are distinct nodes $x$ and $y$ such that every path from $x$ to $y$ traverses $v$ [19]. Since LTP agents are defined for digraphs, each LTP agent in a strongly connected digraph is a strong articulation point; however, the converse does not hold (e.g., node $4$ in Example 2).
The previous discussion describes the properties of the LTP agents. In larger networks, however, multiple LTP agents may exist. This leads to the question: Can a non-influential agent be persuaded by more than one LTP agent?
Lemma 2
Consider a network $\mathcal{G}$ and let $\mathcal{L}$ denote the set of LTP agents in $\mathcal{G}$. Then, $\mathcal{N}_{i}\cap\mathcal{N}_{j}=\emptyset$ for any distinct $i,j\in\mathcal{L}$.
Proof 4.1.
Suppose the contrary is true and there exists a node $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$. Then, by definition, each path in $\mathcal{G}$ from every influential agent in $\mathcal{I}_{q}$ to $q$ must traverse both $i$ and $j$. The following scenarios can occur:
  1. C1
    Every path from $s$ to $q$ traverses $i$ before $j$ for all $s\in\mathcal{I}_{q}$.
  2. C2
    Every path from $s$ to $q$ traverses $j$ before $i$ for all $s\in\mathcal{I}_{q}$.
  3. C3
    Consider $s,u\in\mathcal{I}_{q}$. There exists a path from $s$ to $q$ that traverses $i$ before $j$, and there exists another path from $u$ to $q$ that traverses $j$ before $i$. Here, $s$ and $u$ can be the same node as well.
Under C1, every path from an agent in $\mathcal{I}_{q}$ to $j$ also traverses $i$. Note that $j$ must be a non-influential agent under C1; otherwise $i$ will not be an LTP agent (as a path $j\to q$ exists that does not traverse $i$). Since $j$ has a path to $q$, it follows that $\mathcal{I}_{j}\subseteq\mathcal{I}_{q}$, and hence every path from each node in $\mathcal{I}_{j}$ to $j$ traverses $i$. Thus, $j\in\mathcal{N}_{i}$, so $j$ cannot be an LTP agent. Equivalently, under C2, we obtain $i\in\mathcal{N}_{j}$, so $i$ is not an LTP agent. Hence, both C1 and C2 result in contradiction.
Under C3, there exists a path $P_{1}$ from $s$ to $q$ that traverses $i$ before $j$ and another path $P_{2}$ from $u$ to $q$ that traverses $j$ before $i$. We can represent $P_{1}$ as $s\to i\to j\to q$ and $P_{2}$ as $u\to j\to i\to q$. By definition, a path does not contain repeated nodes. It follows from $P_{1}$ that a path $j\to q$ exists that does not traverse $i$. Further, we know from $P_{2}$ that a path from $u\to j$ does not traverse $i$. Therefore, a path from $u\to q$ exists that does not traverse $i$. Consequently, $q\notin\mathcal{N}_{i}$ which contradicts our assumption that $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$. Similarly, we can show that a path from $s\to q$ exists that does not traverse $j$. Since all of the conditions: C1, C2 and C3, lead to a contradiction, the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.
Lemma 2 establishes that an LTP agent and the agents it persuades form a disjoint group in the network. Certain agents in $\mathcal{G}$ can exist who are neither an LTP agent nor are persuaded by any LTP agent. Excluding them, we can uniquely partition the remaining agents into disjoint groups, with each group consisting of one LTP agent and its persuaded agents.
4.1 The role of network topology on opinion formation
In this subsection, we present a framework that relates the final opinions of agents with the underlying network topology. Subsequently, we employ this framework to demonstrate that the topological relation between an LTP agent and the agent(s) it persuades leads to the formation of an opinion cluster.
Consider a network $\mathcal{G}$ of $n$ agents with opinions evolving under the FJ model (1). Let $\mathcal{G}$ have $m$ stubborn agents. Then, the final opinions satisfy: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$. We can rewrite these linear equations as follows:
$\underbrace{\begin{bmatrix}(I-(I-\beta)W)&-\eta&\\ 0&0&\end{bmatrix}}_{R}\ \underbrace{\begin{bmatrix}\mathbf{x}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}}_{z}=\mathbb{0}$ (2)
where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the $m$ stubborn agents. The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$. Henceforth, we represent eqn. (2) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
Remark 4.2.
The following are the salient properties of $R$: (a) each row sum in $R$ is 0, (b) the diagonal entries are non-negative while the off-diagonal entries are non-positive. Thus, $R$ is a Laplacian matrix under the Defn. 6.3 in [20].
The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$. The associated graph $\mathcal{G}(R)$ has $n+m$ nodes. Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents. Each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$. The source $n+i$ has only one outgoing edge to the node associated with the final opinion of the corresponding stubborn agent. For example, in Fig. 3, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has outgoing edge $(n+1,2)$.
Example 3
Consider the network $\mathcal{G}$ in Fig. 1(a) with stubborn agents $2$ and $6$. The corresponding network $\mathcal{G}(R)$, shown in Fig. 3, is derived from eqn. (2). Each node associated with the initial opinion of a stubborn agent is shown in turquoise.
Figure 3: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
Remark 4.3.
Under Kron Reduction, we determine the Schur complement (4) of the loopy Laplacian matrix $Q$. Since $R$ is a Laplacian matrix, it can be reduced by Kron Reduction for a suitable choice of the subset of nodes $\alpha$. If $\mathcal{G}$ does not have any oblivious agents, then each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$. Thus, by Statement 1) in Lemma 5.1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$.
On the contrary, if $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path in $G(R)$ to this agent. For such networks, the set $\alpha$ must include the oblivious agents belonging to each iSCC and sources $\{n+1,...,n+m\}$. For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$. Thus, Statement 1) in Lemma 5.1 holds.
Consider a set of linear equations $M\mathbf{y}=\mathbb{0}$, where $\mathbf{y}\in\mathbb{R}^{p}$. Eliminating $y[\alpha^{c}]$ using Schur complement (4), yields the reduced system of equations: $M/\alpha^{c}\ \mathbf{y}[\alpha]=0$ (Pg. 18 in [21]). Similarly, we eliminate the states $z[\alpha^{c}]$ in the linear eqns. (2) using the Schur Complement to obtain,
$R/\alpha^{c}.\begin{bmatrix}\mathbf{x}^{*}[\omega]&\mathbf{x}_{s}(0)]\end{bmatrix}^{T}=\mathbb{0}$ (3)
where $\omega=\alpha\setminus\{n+1,...,n+m\}$. Thus, the Kron Reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents.
4.2 Main Results
Using the Kron Reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
Theorem 1.
Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents. Let Assumption 1 hold. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1). If $p$ is an LTP agent, then $p$ and the nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
The proof is given in the Appendix.
In [15], the authors show that the final opinion of an agent depends on the paths (the weights of edges on the path) to it from the stubborn agents. However, it remains unclear what topological conditions on the paths from stubborn agents to distinct agents ensure that the final opinions of these agents converge to the same value. Theorem 1 addresses this gap by proving that the topological relation between an LTP agent and the agents it persuades leads to the formation of an opinion cluster. A major advantage of this result is that regardless of edge weights in the network $\mathcal{G}$ and the stubbornness values, an LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster. Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ and agents in $\mathcal{N}_{3}=\{4,5\}$ form an opinion cluster. Similarly, the LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$.
Remark 4.4.
In [12], the authors derived the conditions to obtain $k\in\mathbb{N}$ opinion clusters in weakly connected digraphs, under the assumption that each cycle in the digraph has at most one stubborn agent. Theorem 1 generalises these conditions to form opinion clusters in any arbitrarily connected digraphs, thereby removing this restrictive assumption.
Moreover, in [12] an opinion cluster is formed by group of agents composed either (1) of a single stubborn agent with the rest being non-stubborn, or (2) of all the oblivious agents in the network $\mathcal{G}$ when $\mathcal{G}$ has a unique aperiodic iSCC of oblivious agents. Hence, a group containing only non-influential agents does not form any opinion cluster in [12]. In contrast, Theorem 1 establishes that an opinion cluster can be formed by any group of non-influential agents if one of them is an LTP agent $p$ and the rest are in $\mathcal{N}_{p}$ (e.g., the agents $3,4,5$ in Example 1).
Remark 4.5.
Consider an agent $p\in\mathcal{V}$. To verify if $p$ is an LTP agent, we first remove all the outgoing edges of $p$, which takes $O(n)$ time. Next, the modified $\mathcal{G}$ is traversed $m+t$ times using the Depth First Search (DFS), once from each of the $m$ stubborn agents and the $t$ iSCCs composed of oblivious agents. As one iteration of DFS has complexity $O(n+|\mathcal{E}|)$, this step requires $O((m+t)(n+|\mathcal{E}|))$ time, where $|\mathcal{E}|$ denotes the number of edges. If a node remains unvisited after all these traversals, then $p$ satisfies Condition (i) in Defn 2. Let $\mathcal{U}_{p}$ be the set of unvisited nodes. By repeating this step for all $p\in\mathcal{V}$, we identify all agents that satisfy Condition (i) and store them as $\mathcal{H}$. This step has a time complexity of $O((t+m)n(n+|\mathcal{E}|))$ which reduces to $O(n(n+|\mathcal{E}|))$ since $t+m\ll n$. If $\mathcal{H}\subseteq\mathcal{I}$, identifying all LTP agents requires $O(n(n+|\mathcal{E}|))$ time.
Next, we verify condition (ii) in Defn. 2 for each non-influential agent in $\mathcal{H}$. For an non-influential agent $p$, we check if there is a $c\in\mathcal{H}\setminus\{p\}$ such that $p\in\mathcal{U}_{c}$. This step requires $O(n^{2})$ time. Repeating it for all $p\in\mathcal{H}$ yields a worst-case complexity of $O(n^{3})$. Thus, the overall time complexity of identifying all LTP agents in $\mathcal{G}$ is $O(n^{3})$.
Using the topological conditions in Theorems 1, a network can be designed to obtain a predefined number of opinion clusters with each cluster formed by a desired set of agents.
Corollary 2.
Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents and $t\geq 0$ iSCCs with oblivious agents such that Assumption 1 holds. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1). Let $\mathcal{O}$ be the set of the oblivious agents that either belong to an iSCC or have paths from only one iSCC in $\mathcal{G}$. Then, $k\geq t+m$ opinion clusters form in $\mathbf{x}^{*}$ if the agents in set $\mathcal{V}\setminus\mathcal{O}$ can be partitioned into ktk-t disjoint groups with each group composed of an LTP agent and the nodes persuaded by it.
The proof of Corollary 2 is given in the Appendix. The following example illustrates the conditions presented in Corollary 2 to obtain the desired opinion clustering.
Example 4
Consider the network $\mathcal{G}$ with stubborn agents $8$ and $10$ in Fig. 4(a). The agents $1,2$ and $3$ are oblivious agents which together form an iSCC (highlighted by a dashed red circle); the remaining agents are non-oblivious. The non-oblivious agents in $\mathcal{G}$ form $4$ disjoint groups highlighted by the green dashed boxes in Fig. 4(a). Each group has an LTP agent (the node with green boundary) and the remaining agents in the group are persuaded by it. For initial conditions from a uniform distribution over $[0,10]$ and stubbornness values in $[0,1)$, the opinions of agents in $\mathcal{G}$ under the FJ model form $5$ opinion clusters as shown in Fig. 4(b). One cluster is formed by the oblivious agent belonging to the iSCC and the rest are formed by the agents belonging to the same group.
(a) Network $\mathcal{G}$
(b) Formation of desired opinion clusters
Figure 4: Suitable design of $\mathcal{G}$ for desired opinion clusters.
5 Conclusions
The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models. Due to varied biases, even the individuals in a closely connected group hold diverse opinions. Even then, a subgroup of individuals can still reach consensus, thereby, forming an opinion cluster. This work examines the role of the underlying network topology in the emergence of these clusters. To begin with, we define a topologically special agent called an LTP agent, and a set of non-influential agents that it persuades. Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent. Using Kron Reduction, we establish in Theorem 1 that an LTP agent and the set of agents it persuades form an opinion cluster in any arbitrarily-connected digraph under the FJ framework. Some additional key insights from our results are as follows: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents. (ii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [12] to any arbitrary digraphs. (iii) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering.
In future, we plan to use the LTP agents for desirably shaping the final opinions in real-world social networks.
Appendix
Kron Reduction: Let $M\in\mathbb{R}^{p\times p}$ with index set $\alpha\subseteq[p]$ and $\alpha^{c}:=[p]\setminus\alpha$. If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
$M/\alpha^{c}=M[\alpha]-M[\alpha,\alpha^{c}](M[\alpha^{c}])^{-1}M[\alpha^{c},\alpha]$ (4)
Under Kron Reduction, the Schur complement of the loopy Laplacian matrix $Q$ is evaluated.
Lemma 5.1 ([18]).
Consider a loopy Laplacian matrix $Q\in\mathbb{R}^{p}$ with $\alpha\subset[p]$. Let $\mathcal{G}(Q)$ be the digraph associated with $Q$.
  1. 1.
    $Q/\alpha^{c}$ is well defined ( i.e. $Q[\alpha^{c}]$ is invertible) if for each node $i\in\alpha^{c}$ there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$. 111(Note that we adapt this condition to our framework because we consider $w_{ij}>0$ when $(j,i)\in\mathcal{E}$.)
  2. 2.
    If $Q$ is a loopless Laplacian matrix, then $Q/\alpha^{c}$ is also a loopless Laplacian matrix.
5.1 Proof of Theorem-1
Before moving to the proof of Theorem-1, the following Lemma needs to be stated.
Lemma 5.2.
Consider a weakly connected network $\mathcal{G}$ with $m\geq 1$ stubborn agents and matrix $R$ derived using eqn. (2). Let $p$ be an LTP agent and $q\in\mathcal{N}_{p}$. If $\alpha$ is the union of set $\{p,q,n+1,....,n+m\}$ and all the oblivious agents in iSCC in $\mathcal{G}$ and $\alpha^{c}=[n+m]\setminus\alpha$ such that $\alpha^{c}\neq\phi$. Then,
$(R[\alpha^{c}])^{-1}=\sum_{k=0}^{\infty}\big((I-\beta[\alpha^{c}])W[\alpha^{c}]\big)^{k}$ (5)
Proof 5.3.
By definition of $\alpha$, we get $\alpha^{c}\subset[n]$. From the construction of $R$, it follows that $R[\alpha^{c}]=I-(I-\beta[\alpha^{c}])W[\alpha^{c}]$. If $\rho((I-\beta[\alpha^{c}])W[\alpha^{c}])<1$, then identity (5) follows by Neumann series. Let $M=(I-\beta[\alpha^{c}])W[\alpha^{c}]$. To prove that $\rho(M)<1$, it suffices to show that (I) $M$ is row substochastic, (II) for each $i^{th}$ row in $M$ with row-sum is $1$ there exists a $j^{th}$ row with row-sum less than $1$ such that there is a path from $j$ to $i$ in the associated digraph $\mathcal{G}(M)$. Then, by Theorem 4.11 in [20])1, $\rho(M)<1$.
Step-I: $M$ is row-substochastic if: (i) a node $i\in\alpha^{c}$ whose in-neighbour is $p,q$ or an oblivious agent in iSCC satisfies $\sum_{j\in\alpha^{c}}w_{ij}<1$, (ii) node $i\in\alpha^{c}$ is a stubborn agent and the corresponding row-sum is at most $(1-\beta_{i})<1$. In an arbitrary digraph $\mathcal{G}$, the following scenarios may occur:
  • $h\in\alpha^{c}$ has an oblivious agent: Due to weak connectivity of $\mathcal{G}$, an oblivious agent in iSCC exists in $\alpha$ that has a path to $h$, implying that (i) holds.
  • All agents in $\alpha^{c}$ are non-oblivious and a stubborn agent exists in $\alpha^{c}$: in this case, (ii) holds for the stubborn agent.
  • All agents in $\alpha^{c}$ are non-oblivious and non-stubborn: Since $q$ cannot be stubborn, node $p$ is the only stubborn agent in $\mathcal{G}$. Thus, at least one node exists in $\alpha^{c}$ that has either $p$ or $q$ as in-neighbour, implying that (i) holds.
Thus, there always exists a node in $\alpha^{c}$ that satisfies either (i) or (ii). Hence, $M$ is row substochastic.
Step-II: As discussed in Step-I, each oblivious agent has a path from an oblivious agent in iSCC in $\mathcal{G}$. It follows from (i) that this path traverses an oblivious agent in $\alpha^{c}$ whose row-sum in $M$ is less than $1$. A special case could occur if each path in $\mathcal{G}$ from the oblivious agents in iSCC to an oblivious agent $\tilde{j}\in\alpha^{c}$ either traverses $p$ or $q$. In this case, the in-neighbour(s) of $p$ or $q$ in $\alpha^{c}$ who has row-sum in $M$ less than $1$ lie on this path. Hence, every oblivious agent in $\alpha^{c}$ with row sum $1$ has a path in $\mathcal{G}(M)$ from a node with row sum <1<1. A similar argument applies to the non-oblivious agents as well. Thus, statement II also holds and $\rho(M)<1$.
Proof of Theorem-1: Let $q\in\mathcal{N}_{p}$. To prove that the final opinion $x_{p}^{*}=x_{q}^{*}$, we reduce the matrix $R$ in eqn. (2) using the Kron Reduction and examine the relation between $x_{p}^{*}$ and $x_{q}^{*}$ in the reduced set of linear equations (3). Let the first $u$ agents in $\mathcal{G}$ be oblivious agents that form the iSCCs. The oblivious agents in each iSCC in $\mathcal{G}$ have consensus [9]. Hence, we consider $q$ is not an oblivious agent in an iSCC; otherwise, $p$ and $q$ belong to an iSCC that is known to achieve consensus.
Consider $\alpha$ to be $\{p,q,1,...,u,n+1,...,n+m\}$. From Remark 4.3, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined. Next, we evaluate $R/\alpha^{c}$ given by eqn. (4). The following structural property of the associated network $\mathcal{G}(R)$ will be used in determining $R/\alpha^{c}$:
  1. P1:
    Each walk that exists from $s\in\mathcal{I}_{q}$ to $q$ in $\mathcal{G}$ traverses $p$ for all $s\in\mathcal{I}_{q}$. Thus, by construction of $\mathcal{G}(R)$, each walk that exists from $j$ to $q$ in $\mathcal{G}(R)$ will traverse $p$ for all $j\in\{[u],n+1,...,n+m\}$.
Property P1 implies that $r_{q,j}=0$ for all $j\in\{[u],n+1,...,n+m\}$ because node $j$ cannot be an in-neighbour of $q$ in $\mathcal{G}(R)$. Additionally, the nodes n+1,,n+mn+1,...,n+m in $\mathcal{G}(R)$ are sources. Thus, the corresponding rows of $R$ have all entries as zero. Therefore, $R[\alpha]$ takes the following form:
$R[\alpha]={\begin{bmatrix}r_{p,p}&r_{p,q}&R[p,[u]]&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&\mathbb{0}&0&...&0&\\ \mathbb{0}&\mathbb{0}&R[[u]]&\mathbb{0}&\cdots&\mathbb{0}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}}$.
Next, we consider $Y=R[\alpha,\alpha^{c}](R[\alpha^{c}])^{-1}R[\alpha^{c},\alpha]$. From Lemma 5.2, it follows that
$Y=R[\alpha,\alpha^{c}]\sum_{k=0}^{\infty}\big((I-\beta[\alpha^{c}])W[\alpha^{c}]\big)^{k}R[\alpha^{c},\alpha]$ (6)
Let $F:[{u}+m+2]\to\alpha$ be a mapping such that $F(1)=p,F(2)=q$, $F(j+2)=j$ for $j\in[u]$, and $F(u+2+j)=n+j$ for $j\in[m]$. This mapping relates the row (and column) indices of $Y$ (and $R/\alpha^{c}$) with the agents in set $\alpha$. Then, from eqn. (6), the entry $y_{ij}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $F(j)$ to $F(i)$ such that each node on this walk (except $F(i)$ and $F(j)$) belongs to $\alpha^{c}$. By property P1), $y_{2j}=0$ for all $j\in{\{3,...,u+m+2\}}$ because each walk from corresponding $F(j)$ to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$. Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the form:
$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&R_{1}&r_{1,u+2}^{1}&...&r_{1,u+m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&\mathbb{0}_{u}&0&...&0\\ \mathbb{0}_{u\times 1}&\mathbb{0}_{u\times 1}&R[[u]]&\mathbb{0}_{u\times 1}&\cdots&\mathbb{0}_{u\times 1}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$}
where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+2]$. The matrix $R_{1}$ is the submatrix $R/\alpha^{c}[1,[u]]$ highlighting the corresponding non-zero entries. From Lemma 5.1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0. Since the second row of $R/\alpha^{c}$ has only two non-zero entries, thus, $r_{2,1}^{1}=-r_{2,2}^{1}$. Hence, by eliminating the states $\alpha^{c}$ in the steady state eqns. (2), we get $x_{p}^{*}=x_{q}^{*}$. Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$. $\blacksquare$
5.2 Proof of Corollary 2
We classified the agents in $\mathcal{G}$ into sets $\mathcal{O}$ and $\mathcal{V}\setminus\mathcal{O}$. As $\mathcal{V}\setminus\mathcal{O}$ is union of tkt-k disjoint sets $\{i\}\cup\mathcal{N}_{i}$, where each $i\in\mathcal{L}$ is an LTP agent. Hence, ktk-t LTP agents are present in set $\mathcal{V}\setminus\mathcal{O}$. By Theorem 1, each LTP agent $i$ and the agents in $\mathcal{N}_{i}$ form an opinion cluster. Therefore, the agents in set $\mathcal{V}\setminus\mathcal{O}$ form ktk-t clusters. The set $\mathcal{V}\setminus\mathcal{O}$ contains the $m$ stubborn agents. Since stubborn agents persuaded by an LTP agent, each stubborn agent must be an LTP agent. Thus, $k-t\geq m$.
[…]
Let $C_{j}$ be the set of oblivious agents in $j^{th}$ iSCC and those oblivious agents only have a path from the $j^{th}$ iSCC. The opinions of agents in $C_{j}$ evolve as $\mathbf{y}_{j}(k)=W_{i}^{j}\mathbf{y}_{j}(k)$ with $\mathbf{y}_{j}$ is the vector that denotes the opinion of agents in this set. Since the associated digraph $G(W_{i}^{t})$ is a subgraph of $\mathcal{G}$ containing a unique aperiodic iSCC, the opinions of the agents converge to consensus (Theorem 5.1 in [20])1.
Hence, the $t$ subgroups, each composed of an iSCC of oblivious agents and those having a path only from the iSCC, form $t$ opinion clusters. Collectively, the oblivious agents in $\mathcal{O}$ and the agents in $\mathcal{V}\setminus\mathcal{O}$ form $k$ opinion clusters such that $k\geq t+m$.